OpenStudy (anonymous):

2 out of 10 people have an allergic reaction to a new medicine. What is the probability that exactly 4 out of 10 people given the medicine experience an allergic reaction?

Parth (parthkohli):

I do not understand the kind of this question :/

OpenStudy (anonymous):

neither do I :( Poop

Parth (parthkohli):

Either, the probability is 0, or half of 2/10 i.e 1/10

Parth (parthkohli):

Is the answer given there?

OpenStudy (anonymous):

Nope.

Parth (parthkohli):

I'm not sure if the question means "4 out of THOSE 10 people" or "4 out of any 10 people"

Parth (parthkohli):

If its 4 out of THOSE 10 people, then 0. If its 4 out of any 10 people, then 1/10

OpenStudy (anonymous):

Are you honestly only twelve?

Parth (parthkohli):

How do you know that? Stalking me eh?

OpenStudy (anonymous):

I looked at your profile, hahah

Parth (parthkohli):

lol looking at anyone's profile = stalking xD

Parth (parthkohli):

You found this from StudyBlue?

OpenStudy (anonymous):

WOAH, hahaha. and found what from what?

Parth (parthkohli):

This question.. where did you get it from?

OpenStudy (anonymous):

A quiz, hhah.

Parth (parthkohli):

And no answer choices? So inconsiderate people heh

OpenStudy (anonymous):

You can not only be 12! I am a highschool student and I mean I am not being tutored by someone 5 years younger than me.

Parth (parthkohli):

Lol all Indians are geeks ;D

Parth (parthkohli):

And it's because of this site that I know all this stuff or I'd have been getting confused Pre-Algebra now

Parth (parthkohli):

confused by^

OpenStudy (anonymous):

On the real though you arent 12.... But anyways thank you for all your help, you seriously saved my "donkey" hahaha.

Parth (parthkohli):

Oh it's question number 17

OpenStudy (anonymous):

What?

Parth (parthkohli):

Just write 1/10.. I'm sure that'd be correct here

Parth (parthkohli):

Inverse relation, y'know

OpenStudy (kropot72):

\[P(4)=\left(\begin{matrix}10 \\ 4\end{matrix}\right)(0.2)^{4}(1-0.2)^{10-4}=\frac{10!}{4!6!}(0.2)^{4}(0.8)^{6}\]This can be solved by the binomial distribution